In the following questions, assume that an adult female is randomly selected. Females have pulse rates that are normally distributed with a mean of 74.0 beats per minute (bpm) and a standard deviation of 12. 5 bpm. Round your answers to 4 decimal places.

a. Find the probability of a pulse rate less than 100 bpm

b. Find the probability of a pulse rate greater than 80 bpm

c. Find the probability of a pulse rate between 60 and 70 bpm

d. Find the probability of a pulse rate between 70 bpm and 90 bpm
Answer #1

Given that, el = 74.0 bpm 6= 12.5 bpm 0x<100)=P(x-el 100-74 12.5 =P(Z < 2.08) F. P(x <100) = 0.9812b1 P(x > 80) 1 - P(x580) as p(x-4 < 80-74) - 1 p (z 40.48) - 1-0.6844 HP(x780) = 0.3156 CP ( 6o <x (70) - pr 60-74 12 .5 X- 7​​​​​​업 P(70CxC99 = PCIe < 0) - IP (-0.32 / 2 - 1.28) - P ( x 1.28) - P(x--6.32) = 0-8997 - 0.3745 -0.5252 -P70<x/goj - 0:5252